the sum of two numbers is 172. the first is 8 less than 5 times the second. find the fist number.
Answer by checkley77(12844)

X+Y=172 OE X=172-Y
X=5Y-8
5Y-8=172-Y
5Y+Y=172+8
6Y=180
Y=180/6
Y=30 ANS
X=172-30
X=142 ANS. FOR THE FIRST NUMBER WHICH IS:
PROOF:
142=5*30-8
142=150-8
142=142